OpenStudy (anonymous):

If cos Θ = square root 2 over 2 and 3 pi over 2 < Θ < 2π, what are the values of sin Θ and tan Θ?

OpenStudy (anonymous):

Can u just type it as equation

OpenStudy (anonymous):

@jamieallover

OpenStudy (anonymous):

@jamieallover22

OpenStudy (anonymous):

Yeah, sorry. Here's a screenshot.

OpenStudy (anonymous):

Like I know what the actual answer is, I just don't know how to get it.

OpenStudy (anonymous):

@sheelo_mughal

OpenStudy (anonymous):

draw a diagram

OpenStudy (anonymous):

Did the screenshot not load? @sourwing

OpenStudy (anonymous):

If I drew it you wouldn't be able to tell what it was.

OpenStudy (anonymous):

It loaded, which quadrant is theta in?

OpenStudy (anonymous):

It doesn't say, but if I look at my trig graph paper it looks like it's in either I or IV cause it's between 3pi/2 and 2pi

OpenStudy (anonymous):

between 3pi/2 and 2pi is the IV quadrant. 

OpenStudy (anonymous):

what is the length of the missing side?

OpenStudy (anonymous):

Is it \[\sqrt{2}\] ??

OpenStudy (anonymous):

That's what I get if I solve a^2+b^2=c^2

OpenStudy (radar):

What is the sign of the tangent function in quadrant IV? What is the sign of the sine function in quadrant IV? If you know that you can solve by inspection.

OpenStudy (anonymous):

it's -sqrt(2) since theta is in the 4th quadrant